The question says: solve each equation by completing the square.
m to the second power +14m-58=-7.

Found 2 solutions by checkley71, ankor@dixie-net.com:
Answer by checkley71(8403)

M^2+14M-58=-7
M^2+14M=-7+58
M^2+14M=51
M^2+14M+49=51+49
(M+7)(M+7)=100
M+7=10
M=10-7
M=3
PROOF
3*3+14*3-58=-7
9+42-58=-7
51-58=-7
-7=-7

Answer by ankor@dixie-net.com(12704)

The question says: solve each equation by completing the square.
m to the second power +14m-58=-7.
:
m^2 + 14m - 58 = - 7
:
m^2 + 14m + __ = -7 + 58
:
Find the value to complete the square (14/2)^2 = 49; add 49 to both sides
m^2 + 14m + 49 = 51 + 49
m^2 + 14m + 49 = 100
(m+7)^2 = 100
:
Find the square root of both sides:
x + 7 = +/-SqRt(100)
x = -7 + 10
x = 3
and
x = -7 - 10
x = -17
:
x = 3 and x = -17
: